OpenStudy (anonymous):

Evaluate the double integral of (e^-(x^2+(y-x)^2+y^2))dA with the bounds -inf

OpenStudy (anonymous):

To clarify: the integral is \[\large\int_{-\infty}^\infty\int_{-\infty}^\infty \exp\bigg[-\left(x^2+(y-x)^2+y^2\right)\bigg]~dx~dy~~?\]

OpenStudy (anonymous):

You must be away or something. I'll assume the integral is as described above. The region \(x\in(-\infty,\infty),y\in(-\infty,\infty)\) is the entire \(x-y\) plane, so when transforming to polar coordinates, you have \(r\in(0,\infty)\) and \(\theta\in(0,2\pi)\). In polar coordinates, you have \[\begin{cases}x=r\cos\theta\\y=r\sin\theta\\dx~dy=r~dr~d\theta\end{cases}\] So the integral itself, post-transformation, is equivalent to \[\large\int_0^{2\pi}\int_0^\infty\exp\bigg[-\left(r^2+(r\sin\theta-r\cos\theta)^2\right)\bigg]~r~dr~d\theta\\ \large\int_0^{2\pi}\int_0^\infty\exp\bigg[-r^2-r^2(\sin\theta-\cos\theta)^2\bigg]~r~dr~d\theta\] Can you take it from here?

OpenStudy (anonymous):

Yeah, thanks. For some reason, the professor recommended we utilize a jacobian, but I just don't see how it's practical with this problem. I guess I'll find out later. Again, thanks for the help.

OpenStudy (anonymous):

Well the Jacobian is still used here. \[f(r,\theta)=r\cos\theta\\ g(r,\theta)=r\sin\theta\] \[J=\begin{vmatrix}\frac{\partial f}{\partial r}&\frac{\partial f}{\partial\theta}\\ \frac{\partial g}{\partial r}&\frac{\partial g}{\partial\theta}\end{vmatrix}=\begin{vmatrix}\cos\theta&-r\sin\theta\\ \sin\theta&r\cos\theta\end{vmatrix}=r\cos^2\theta+r\sin^2\theta=r\] Generally, the Jacobian is used in any change of variables, and for the polar conversion you get: \[\int\int_Rf(x,y)~dx~dy~~\iff~~\int\int_Sf(r,\theta)J~dr~d\theta\] where \(J=r\).